I am completely lost and I can't seem to find the right answer...
(3,-6);3x=5y=4
Write an equation of the line containing the given point and perpindicular to the line
y=mx=b

Found 2 solutions by Alan3354, stanbon:
Answer by Alan3354(69443)

(3,-6);
3x=5y=4 too many equal signs

Answer by stanbon(75887)

(3,-6);3x+5y=4
Write an equation of the line containing the given point and perpindicular to the line
-----
Find the slope of the given line:
5y = -3x+4
y = (-3/5)x + (4/5)
----
slope = -3/5
----
So, find the equation of the line with
slope = (-3/5) and y = -6 when x = 3.
----
Equation:
y = mx + b
-6 = (-3/5)*3 + b
-6 = -9/5 + b
-30/5 = -9/5 + b
b = -21/5
--------------
y = (-3/5)x - (21/5)
=======================
Cheers,
Stan H.
==============

----------------
Cheers,
Stan H.
==========